OpenStudy (anonymous):

abs. sin theta -tan theta/tan theta=0.1

OpenStudy (anonymous):

\[\left| \sin \theta-\tan \theta \right|/\tan \theta=0.1\]

OpenStudy (anonymous):

what is theta in degrees?

OpenStudy (anonymous):

Only the numerator is in the absolute value bars?

OpenStudy (rsadhvika):

\(\large \frac{\left| \sin \theta-\tan \theta \right|}{\tan \theta}=0.1 \)

OpenStudy (rsadhvika):

split the equation into two, and solve it as u wud solve any absolute value equation

OpenStudy (rsadhvika):

\(\large \left| \sin \theta-\tan \theta \right| = 0.1\tan \theta \) \(\large \sin \theta-\tan \theta = 0.1\tan \theta \) OR \(\large \sin \theta-\tan \theta = -0.1\tan \theta \)

OpenStudy (anonymous):

i keep getting \[\cos \theta =1.1\]

OpenStudy (anonymous):

which is not in the domain

OpenStudy (rsadhvika):

that means you're doing it correctly :) simply discard that value